OpenStudy (twizttiez):

What is the value of the function y=2x+3 when x = –1? A. 5 B. –5 C. 2 D. 1

OpenStudy (solomonzelman):

\(\large\color{slate}{ y=2x+3 }\) ~~~~~~~~~~~~~~~~~~~~~~~~~ when x=3, then \(\large\color{slate}{ y=2x+3 }\) \(\large\color{slate}{ y=2(3)+3 }\) \(\large\color{slate}{ y=6+3 }\) \(\large\color{blue}{ y=9 }\) ~~~~~~~~~~~~~~~~~~~~~~~~~ when x=0, then \(\large\color{slate}{ y=2x+3 }\) \(\large\color{slate}{ y=2(0)+3 }\) \(\large\color{slate}{ y=0+3 }\) \(\large\color{blue}{ y=3 }\) ~~~~~~~~~~~~~~~~~~~~~~~~~ when x=-2, then \(\large\color{slate}{ y=2x+3 }\) \(\large\color{slate}{ y=2(-2)+3 }\) \(\large\color{slate}{ y=-4+3 }\) \(\large\color{blue}{ y=-1 }\) can you tell me what is it, when x=-1?